OpenStudy (anonymous):

A parent can choose from 6 types of protein, 5 vegetables and 7 desserts. If the parent serves a meal of 1 protein, 1 vegetable and 1 dessert to the family, how many different meals can be served?

OpenStudy (anonymous):

6*5*7

OpenStudy (anonymous):

1 protein out of 6 can be chosen in : \(^6C_1\) 1 vegetable our of 5 can be chosen in: \(^5C_1\) 1 dessert out of 7 can be chosen in : \(^7C_1\) So: \[^6C_1 \times ^5C_1 \times ^7C_1 \implies 6 \times 5 \times 7 = ??\]